### v4

@@ 4.1 Set-theoretic representation of Forman-Ricci curvature
Algorithm 3 Compute Forman-Ricci Curvature: Method B
  Input: $C,d_{\max}\geq 1$
  $\mathcal{F}_{d}=\emptyset,\forall d\leq d_{\max}$
  Compute Neigh(G) from Algorithm 1;
  for $\alpha\in C$ do
     $d:=|\alpha|-1$
     Compute $\mathrm{F}_{d}(c)$ according to (21):
     Compute $\pi_{\alpha};$
     Sset $p:=0;$
     for  $\gamma\in\partial(\alpha)$ do
        set $M:=\{v\in V\,|\,v\in\pi_{\gamma},\,v\notin\alpha\}$
        for  $v\in M$ do
           if  $v\notin\pi_{\alpha}$ then
              $p:=p+1$
           end if
        end for
     end for
     set $\mathrm{F}_{d}(\alpha)=|\pi_{\alpha}|+(d+1)-p$
     Update $\mathcal{F}_{d}:=\mathcal{F}_{d}\cup\{\mathrm{F}_{d}(\alpha)\}$
  end for
  return  $\mathcal{F}=\bigcup\limits_{d}\{\mathcal{F}_{d}\}$
Algorithm 4 Compute Forman-Ricci Curvature: Method C
  Input: $C,d_{\max}\geq 1$
  $\mathcal{F}_{d}=\emptyset,\forall d\leq d_{\max}$
  Compute Neigh(G) from Algorithm 1;
  for $\alpha\in C$ do
     $d:=|\alpha|-1$
     Compute $\mathrm{F}_{d}(\alpha)$ according to (24):
     Set $h:=|\pi_{\alpha}|$ and $n:=0;$
     for $\gamma\in\partial(\alpha)$ do
        Update $n:=n+|\pi_{\gamma}|;$
        Set $\mathrm{F}_{d}(\alpha)=(d+2)h+2(d+1)-n;$
     end for
     Update $\mathcal{F}_{d}:=\mathcal{F}_{d}\cup\{\mathrm{F}_{d}(\alpha)\}$
  end for
  return  $\mathcal{F}=\bigcup\limits_{d}\{\mathcal{F}_{d}\}$
Figure 2: Example of how the algorithm detects cell neighbours from node neighbourhood for $2$-cells (triangles). The $2$-cell $\alpha$ is enhanced in orange, whilst its boundary is drawn in red. According to the algorithm, the number of $3$-cells containing $\alpha$ in its boundary ($|H_{\alpha}|$) coincides with the number of common neighbours of the nodes in $\alpha$ ($|\pi_{\alpha}|$). Also, the number of neighbours per boundary is counted by the size of intersections of node neighbourhoods in the boundary. In this example, $\mathrm{F}_{2}(\alpha)=4\cdot|\{6\}|+3-(|\{3,6\}|+|\{6\}|+|\{6\}|)=3.$ The same output is reached by using (5): $\mathrm{F}_{2}(\alpha)=1+3-1=3.$
4.2 Computation and Benchmark of Novel Algorithms
@@
To validate our set-theoretic representation of FRC, we developed a Python language version [24] of our pseudo-codes, which we named FastForman - Method A, B and C, which uses the formulation (19), (21) and (24), respectively. Furthermore, we run benchmark tests and compared FastForman’s execution time and memory usage with the leading software in the literature, namely Python packages HodgeLaplacians [23] and GeneralisedFormanRicci [27]. The benchmark tests were performed on a Dell laptop model XPS 15-7590, Intel Core i7-9750H CPU 2.60GHz, 32Gb RAM, running Linux Ubuntu operating system version 20.04.4 LTS (Focal Fossa). To carry out the tests, with the help of NetworkX package [13], we generated $20$ copies of $3$-dimensional point cloud data from random geometric networks [19] with the parameters $n=50$ nodes and radius $r\in\{0.1,0.2,\ldots,1.0\}$ (to access this data we refer the reader to the link [5]). To undertake an unbiased benchmark test, it is first worth noting that the Python package, GeneralisedFormanRicci, is limited to computing the FRC for edges and triangles. Therefore, we divided the benchmark tests into two separate groups. The first group aimed at comparing GeneralisedFormanRicci, HodgeLaplacians and FastForman (A, B and C) for $d_{max}=2$. The second group focused on comparing HodgeLaplacians versus FastForman (A, B and C) for $d_{max}=5$. To compare the time complexity between all involved software packages, we considered the total execution time from the generation of cells (cliques) to the end computation of FRC. For a fair comparison, we employed the same software package, Gudhi [15], to list the cells of the aforementioned generated networks. Due to the separate benchmark groups, we first depict in Fig. 3 the average number of cells for $d_{\max}\in\{2,5\}$ contained within the aforementioned generated networks. In Fig. 4 we benchmark the average of the total processing time and average memory peak usage for computing the FRC over the generated random networks. We also compare the results with the time and memory consumption of listing the set of cells via the Gudhi package (light blue curves in the figure). We subsequently discuss the merits and limitations of each software package for computing FRC.
@@
GeneralisedFormanRicci: Due to computational limitations with memory management, we could not compute FRC for radius values higher than $r=0.7$ via GeneralisedFormanRicci. Despite GeneralisedFormanRicci being limited to lower dimensional computations, the time consumption was the highest, exceeding the Hodgelaplacians and FastForman implementations. Moreover, it displays a high variance of time and memory consumption particularly for dense networks, which might be associated with its limitations in computing FRC up to 2-dimensional cells (compare Figs. 3 and 4). The limitation to point cloud data sets is also a drawback of the algorithm since it restricts one from selecting a subset of cells over which one may compute FRC.
@@
Hodgelaplacians: The time processing is comparable with the performance results of FastForman for low-dimensional cells and low connectivity values. However, the computational efficiency decreases for high-order cells and larger radii and thus its performance falls behind in comparison to FastForman. Furthermore, the time and memory processing variance increase significantly with respect to the network complexity (ie., size, number of higher-order cells and density). This fact might be explained by the computation of FRC via Bochner’s discrete formula, which uses Hodge Laplacian and Bochner Laplacian matrix operators [2, 26, 4]. This demands memory for storing the matrices and time to both generate the matrices and compute FRC. It is worth emphasising that the Hodge Laplacian matrix for the highest cell dimension is simply the $d_{\max}$-th boundary operator, which in computational terms, induces a miscalculation on FRC for this dimension. Thus, for the correct computation for all cells, the entire simplicial complex should be provided, which impacts directly on the computational complexity. One way to improve efficiency and avoid miscalculation would be to include information of $(d_{\max}+1)$-cells so the correct computation up to $d_{\max}$ can be computed correctly. A detailed report of this special case is included in the link [5], where we provide an example of the miscalculation.
@@
FastForman: Our implementations had equivalent performance and superseded existing FRC software packages benchmarked in our tests. The key to this performance is that our set-theoretic formulation leads us to algorithms that only require the use of local node neighbourhood information for computing the FRC by cell, which, as a consequence, reduces the complexity of computing FRC. Noteworthy, there is no requirement for specifically computing and storing neighbouring cells, hence this explains its low memory usage and fast computation. Moreover, FastForman implementations are versatile in the sense that they do not make mandatory the inclusion of the global information of an entire simplicial complex. In other words, the FRC computation is flexible, in a way that it can be computed for subsets of cells within a given simplicial complex. The limiting factors are twofold: Firstly, the complexity of the algorithms is dependent upon how cells are computed, which is generically an NP-problem and herein we employed the Gudhi algorithm. Secondly, the algorithm’s complexity is associated with node neighbourhood computation, as given by Algorithm 1, which has a time complexity of $\mathcal{O}(|E|)$ since it iterates over edges. Indeed, FastForman’s implementation processing time increases in accordance with the time required for finding cells; see Figs. 4 and 3. To be more precise, the space complexity is, in the worse case (when the graph is complete), $\mathcal{O}(\mathrm{C}(d_{\max})+|V|^{2})$, where $\mathrm{C}(d_{\max})$ is the space complexity for the finding cliques algorithm (up to dimension $d_{\max}$) and $|V|^{2}$ is the space complexity for storing the neighbourhood of nodes, in the worse scenario. The time complexity is $\mathcal{O}(\sum_{d=1}^{d_{\max}}|\mathrm{C}_{d}|\cdot(d+1)\cdot|V|^{(d+1))}$, where $(d+1)$ is the time complexity for iterating the boundary of a $d$-cell and $|V|^{(d+1)}$ is the order complexity for computing set intersection of $(d+1)$ node neighbourhoods (taking into account that we can compute set intersection in $\mathcal{O}(\sum_{x\in\alpha}|\pi_{x}|$) time. As an application, we showcase our optimal performance in random graphs, aiming to extend the results on geometry detection from the work in [26]. From Theorem 7.21, we are able to compare the exact range of values reached by FRC as a function of the number of nodes and the maximum cell dimensions. Moreover, we show we can easily compute the frequency (or distribution) of FRC values in function of the dimension of the cells and compare them across different geometric graphs. To this end, we generated the expected frequency of FRC values for $1000$ random graphs in each of the following cases: Erdős-Renyi, 2D Random Geometric, Barabási-Albert and Watts-Strogatz model (with edge randomization parameter $p=0.25$). All graphs were set up with $n=50$ nodes and an approximate density of $0.5$. We computed the FRC distribuitions for $d\in\{1,2,3,4\}$, which can be seen in Fig. 5. Observe that the standard deviation of the FRC values decreases as the FRC dimension increases. Also, the Watts Strogatz curve becomes less distinguishable from Erdős-Renyi, which may indicate that the random features of Watts Strogatz can be detected only via higher-order analysis. In general, all FRC could be used for geometry detection.
@@
Figure 4: Benchmark for computing Forman-Ricci curvature by using HodgeLaplacians, GeneralisedFormanRicci and FastForman With methods A, B and C, in comparison with the cells computation by using the Gudhi algorithm, for $n=50$ and $d_{\max}\in\{2,5\}$. We display the result in a log scale to clarify the analysis.
Figure 5: Comparison of FRC for higher order dimensions between different random graph models.
@@ 5 Conclusion
We proposed novel set theoretical formulations for high-order Forman-Ricci curvature (FRC) and this led us to the implementations of new algorithms, which we coined FastForman (Method A, B and C). Central to our set-theoretic representation theory for high-order network cells and FRC computation is the fact we managed to redefine parallelism and transversality of neighbouring cells. This enabled us to formulate FRC computation in terms of the neighbours of nodes in a network. As a consequence, we managed to computationally tame some of the combinatorial complexities intrinsic to high-order networks. We validated FastForman algorithms via benchmark tests on random higher-order geometric graphs and compared them with leading softwares, namely, Hodgelaplacians and GeneralisedFormanRicci. To ensure a fair comparison, we implemented FastForman with Python language since Hodgelaplacians and GeneralisedFormanRicci are also implemented in Phyton. We also took into consideration the intrinsic limitations of Hodgelaplacians and GeneralisedFormanRicci, such as the maximum number of cell dimensions tolerated by these algorithms. Our benchmark tests established that FastForman implementations are more efficient in time and memory complexity. In conclusion, despite the constraints of cell search NP algorithms and the combinatorial explosion of high-order network structures, our findings facilitate the computation of higher-order geometric invariants in complex networks. We envisage that FastForman will open novel research avenues in big data science, particularly through a geometrical point of view.
Acknowledgments
The authors would like to acknowledge Rodrigo A. Moreira at the Basque Center for Applied Mathematics for his critical review. This research is supported by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033. Moreover, the authors acknowledge the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. SR further acknowledges the RTI2018-093860-B-C21 funded by (AEI/FEDER, UE) and the acronym “MathNEURO”.
Supplementary Materials
6 Examples
@@
Figure 6: Neighborhood condition depicted, for $d\in\{1,2,3\}$. For all dimensions, the neighbouring cells $\alpha$ and $\alpha^{\prime}$ are drawn in blue and yellow, respectively, whilst the common boundary ($\alpha\cap\alpha^{\prime}$) is represented in green. The presence (or absence) of the red dashed edge in the simplicial complex defines whether $\alpha$ and $\alpha^{\prime}$ are transverse or parallel. In all cases, the intersection $\gamma:=\alpha\cap\alpha^{\prime}$ is a $(d-1)$-cell such that $\gamma<\alpha,\alpha^{\prime}$, which guarantees that for $\alpha$ and $\alpha^{\prime}$ to be neighbours, it is sufficient to reach the cell neighbourhood condition $1$ is reached. This result is a consequence of Theorem 7.7.
@@
In Fig. 6, we have the characterization of neighborhood of $d$-cells, $\alpha$ (blue) and $\alpha^{\prime}$ (yellow), for $d\in\{1,2,3\}.$ Suppose that the red-dashed edge belongs to the graph. For $d=1$, we have that $\alpha=\{2,3\}$ and $\alpha^{\prime}=\{2,4\}$ are transverse neighbors, once that the $2$-cell $\beta=\{2,3,4\}$ is such that $\beta>\alpha,\alpha^{\prime}$. Analogously, for $d=2$, the $2$-cells $\alpha=\{1,2,4\}$ and $\alpha^{\prime}=\{2,3,4\}$ are transverse once that $\beta=\{1,2,3,4\}$ is a $3$-cell such that $\beta>\alpha,\alpha^{\prime}$. Also, for $d=3$, $\alpha=\{1,2,3,4\}$, $\alpha^{\prime}=\{2,3,4,5\}$ are transverse once that $\beta=\{1,2,3,4,5\}$ is a $4$-cell such that $\beta>\alpha,\alpha^{\prime}$. Regardless of the existence of the red-dashed edge in the graph, we have that, for $d\in\{1,2,3\}$, there is a $(d-1)$-cell, $\gamma:=\alpha\cap\alpha^{\prime}$ such that $\gamma<\alpha,\alpha^{\prime}$. This fact can be justified by the result of Theorem 7.3, which guarantees that is sufficient to have a common boundary between two $d$-cells so they can be neighbours. The dashed-red edge’s absence (or presence) guarantees they are parallel (transverse) neighbours.
@@
7 Demonstrations
Here we are considering $\pi_{\alpha}$, $\pi_{x}$ and the concepts of abstract simplicial complexes as defined in Section 2.
Theorem 7.1.

### v5

@@ 4.1 Set-theoretic representation of Forman-Ricci curvature
Note that Algorithm 2 is versatile in the sense that it can compute FRC for an arbitrary subset of cells, not only for the full simplicial complex (provided the correct graph $G$ associated with this simplicial complex, i.e., the graph that generates the provided nodes). Indeed, it can compute FRC cell by cell, and in doing so it only accesses local information at the time at which that cell is being called (i.e. node neighbourhood for each node in the cell). Since it iterates through the set of cells (as parameterized by some cliques’ algorithm) and because it does not store a full simplicial complex, then it results in an algorithm with efficient time and memory footprint. Indeed, the key to this computational efficiency is that it is solely based on local computation thus it significantly facilitates memory management and processing time. The main limiting factor of our algorithm is dependent on the specific cliques’ algorithm used, which is in general NP-complete and is unavoidable to any algorithm with this purpose. One way to reduce the time complexity is by limiting the search space for cliques, such as, by setting an upper bound for the highest dimension of cells ($d_{\max}$), and also using efficient algorithms for clique’s computing.
4.2 Ricci-curvature Computation and Benchmark of Novel Algorithm
@@
To validate our set-theoretic representation of FRC, we developed a Python language version [python] of our pseudo-code (Algorithm 2), which we named FastForman. Furthermore, we run benchmark tests and compared FastForman’s execution time and memory usage with the leading software in the literature, namely Python packages HodgeLaplacians [hodgelaplacians] and GeneralisedFormanRicci [generalisedformanricci]. The benchmark tests were performed on a Dell laptop model XPS 15-7590, Intel Core i7-9750H CPU 2.60GHz, 32Gb RAM, running Linux Ubuntu operating system version 20.04.4 LTS (Focal Fossa). To carry out the tests, with the help of NetworkX package [hagberg2020networkx], we generated $20$ copies of $3$-dimensional point cloud data from random geometric networks [penrose2003random] with the parameters $n=50$ nodes and radius $r\in\{0.1,0.2,\ldots,1.0\}$ (to access this data we refer the reader to the link [kaggle_FRC_report]). To undertake an unbiased benchmark test, it is first worth noting that the Python package, GeneralisedFormanRicci, is limited to computing the FRC for edges and triangles. Therefore, we divided the benchmark tests into two separate groups. The first group aimed at comparing GeneralisedFormanRicci, HodgeLaplacians and FastForman for $d_{max}=2$. The second group focused on comparing HodgeLaplacians versus FastForman for $d_{max}=5$. To compare the time complexity between all involved software packages, we considered the total execution time from the generation of cells (cliques) to the end computation of FRC. For a fair comparison, we employed the same software package, Gudhi [maria2014gudhi], to list the cells of the aforementioned generated networks. Due to the separate benchmark groups, we first depict in Fig. 3 the average number of cells for $d_{\max}\in\{2,5\}$ contained within the aforementioned generated networks. In Figs. 4 and 5 we benchmark the average of the total processing time and average memory peak usage for computing the FRC over the generated random networks. We also compare the results with the time and memory consumption of listing the set of cells via the Gudhi package (grey curves in the figures). We subsequently discuss the merits and limitations of each software package for computing FRC.
@@
GeneralisedFormanRicci: Due to computational limitations with memory management, we could not compute FRC for radius values higher than $r=0.7$ via GeneralisedFormanRicci. Despite GeneralisedFormanRicci being limited to lower dimensional computations, the time consumption was the highest, exceeding the Hodgelaplacians and FastForman implementations. Moreover, it displays a high variance of time and memory consumption particularly for dense networks, which might be associated with its limitations in computing FRC up-to 2-dimensional cells (compare Figs. 3 and 4). The limitation to point cloud data sets is also a drawback of the algorithm since it restricts one from selecting a subset of cells over which one may compute FRC.
@@
Hodgelaplacians: The time processing is comparable with the performance results of FastForman for low-dimensional cells and low connectivity values. However, the computational efficiency decreases for high-order cells and larger radii and thus its performance falls behind in comparison to FastForman. Furthermore, the time and memory processing variance increase significantly with respect to the network complexity (ie., size, number of higher-order cells and density). This fact might be explained by the computation of FRC via Bochner’s discrete formula, which uses Hodge Laplacian and Bochner Laplacian matrix operators [borner2007network, weber2017characterizing, dakurah2022modelling]. This demands memory for storing the matrices and time to both generate the matrices and compute FRC. It is worth emphasising that the Hodge Laplacian matrix for the highest cell dimension is simply the $d_{\max}$-th boundary operator, which in computational terms, induces a miscalculation on FRC for this dimension. Thus, for the correct computation for all cells, the entire simplicial complex should be provided, which impacts directly on the computational complexity. One way to improve efficiency and avoid miscalculation would be to include information of $(d_{\max}+1)$-cells so the correct computation up to $d_{\max}$ can be computed correctly. A detailed report of this special case is included in the link [kaggle_FRC_report], where we provide an example of the miscalculation.
@@
FastForman: Our implementation supersedes existing FRC software packages benchmarked in our tests. The key to this performance is that our set-theoretic formulation leads us to an algorithm that only requires the use of local node neighbourhood information for computing the FRC by cell, which, as a consequence, reduces the complexity of computing FRC. Noteworthy, there is no requirement for specifically computing and storing neighbouring cells, hence this explains its low memory usage and fast computation. Moreover, FastForman is versatile in the sense that it does not make mandatory the inclusion of the global information of an entire simplicial complex. In other words, the FRC computation is flexible, in a way that it can be computed for subsets of cells within a given simplicial complex. The limiting factors are two folds: Firstly, the complexity of the algorithm is dependent upon how cells are computed, which is generically an NP-problem and herein in we employed the Gudhi algorithm. Secondly, the algorithm’s complexity is associated with node neighbourhood computation, as given by Algorithm 1, which has a time complexity of $\mathcal{O}(|E|)$ since it iterates over edges. Indeed, FastForman’s processing time increases in accordance with the time required for finding cells; see Figs. 4 and 3.
To conclude, we provide a brief benchmark summary in Table 1, where we compare FastForman, Hodgelaplacians and GeneralisedFormanRicci. The table summarises the versatility of the algorithms in choosing the set of cells, the ability in computing FRC for higher-order dimensional cells, as well as processing time and memory usage.
Table 1: Algorithm comparison summary
Algorithm Flexible High-order Time processing Memory Peak
GeneralisedFormanRicci No $d_{\max}=2$ Very high Very high
HodgeLaplacians No Yes High High
FastForman Yes Yes Fair Low
@@
Figure 4: Benchmark for computing Forman-Ricci curvature by using HodgeLaplacians (blue), FastForman (red) and GeneralisedFormanRicci (green), in comparison with the cells computation by using the Gudhi algorithm, for $n=50$ and $d_{\max}\in\{2,5\}$.
Figure 5: Benchmark for computing Forman-Ricci curvature by using HodgeLaplacians (blue), FastForman (red) and GeneralisedFormanRicci (green), in comparison with the cells computation by using the Gudhi algorithm, for $n=50$ and $d_{\max}\in\{2,5\}$. Here, we display the result in a log scale to clarify the analysis.
@@ 5 Conclusion
We proposed a novel set theoretical formulation for high-order Forman-Ricci curvature (FRC) and this led us to the implementation of a new algorithm, which we coined FastForman. Central to our set-theoretic representation theory for high-order network cells and FRC computation is the fact we managed to redefine parallelism and transversality of neighbouring cells. This enabled us to formulate FRC computation in terms of the neighbours of nodes in a network. As a consequence, we managed to computationally tame some of the combinatorial complexities intrinsic to high-order networks. We validated FastForman via benchmark tests on random higher-order geometric graphs and compared them with leading softwares, namely, Hodgelaplacians and GeneralisedFormanRicci. To ensure a fair comparison, we implemented FastForman with Python language since Hodgelaplacians and GeneralisedFormanRicci are also implemented in Phyton. We also took into consideration the intrinsic limitations of Hodgelaplacians and GeneralisedFormanRicci, such as the maximum number of cell dimensions tolerated by these algorithms. Our benchmark tests established that FastForman is more efficient in time and memory complexity. In conclusion, despite the constraints of cell search NP algorithms and the combinatorial explosion of high-order network structures, our findings facilitate the computation of higher-order geometric invariants in complex networks. We envisage that FastForman will open novel research avenues in big data science, particularly through a geometrical point of view.
Appendix A Examples
@@
Figure 6: Neighborhood condition depicted, for $d\in\{1,2,3\}$. For all dimensions, the neighbouring cells $\alpha$ and $\alpha^{\prime}$ are drawn in blue and yellow, respectively, whilst the common boundary ($\alpha\cap\alpha^{\prime}$) is represented in green. The presence (or absence) of the red dashed edge in the simplicial complex defines whether $\alpha$ and $\alpha^{\prime}$ are transverse or parallel. In all cases, the intersection $\gamma:=\alpha\cap\alpha^{\prime}$ is a $(d-1)$-cell such that $\gamma<\alpha,\alpha^{\prime}$, which guarantees that for $\alpha$ and $\alpha^{\prime}$ to be neighbours, it is sufficient to reach the cell neighbourhood condition $1$ is reached. This result is a consequence of Theorem B.7.
@@
In Fig. 6, we have the characterization of neighborhood of $d$-cells, $\alpha$ (blue) and $\alpha^{\prime}$ (yellow), for $d\in\{1,2,3\}.$ Suppose that the red-dashed edge belongs to the graph. For $d=1$, we have that $\alpha=\{2,3\}$ and $\alpha^{\prime}=\{2,4\}$ are transverse neighbors, once that the $2$-cell $\beta=\{2,3,4\}$ is such that $\beta>\alpha,\alpha^{\prime}$. Analogously, for $d=2$, the $2$-cells $\alpha=\{1,2,4\}$ and $\alpha^{\prime}=\{2,3,4\}$ are transverse once that $\beta=\{1,2,3,4\}$ is a $3$-cell such that $\beta>\alpha,\alpha^{\prime}$. Also, for $d=3$, $\alpha=\{1,2,3,4\}$, $\alpha^{\prime}=\{2,3,4,5\}$ are transverse once that $\beta=\{1,2,3,4,5\}$ is a $4$-cell such that $\beta>\alpha,\alpha^{\prime}$. Regardless of the existence of the red-dashed edge in the graph, we have that, for $d\in\{1,2,3\}$, there is a $(d-1)$-cell, $\gamma:=\alpha\cap\alpha^{\prime}$ such that $\gamma<\alpha,\alpha^{\prime}$. This fact can be justified by the result of Theorem B.3, which guarantees that is sufficient to have a common boundary between two $d$-cells so they can be neighbours. The dashed-red edge’s absence (or presence) guarantees they are parallel (transverse) neighbours.
@@
Appendix B Demonstrations
Here we are considering $\pi_{\alpha}$ and $\pi_{x}$ as defined in Section 2.
Theorem B.1.